OpenStudy (anonymous):

Use the formula d = rt to find r when d = 200 feet and t = 4 seconds.

OpenStudy (saifoo.khan):

When you have the formula and everything.. then where are you stuck?

OpenStudy (kropot72):

d = rt r = d/t = 200/4 = 50 feet per second